Status Report

NOAA SATOPS Morning Report: Friday, October 12, 2007

By SpaceRef Editor
October 12, 2007
Filed under , , ,

Customers or users who want to call to ask about problems or report outages can use the following contacts:

– For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444

– To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: [email protected]

J285

GOES

Note: Fall eclipse season for GOES. For more information on canceled images due to eclipse season please see: http://www.oso.noaa.gov/goes/koz-eclipse/koz-eclipse.htm

All GOES operations were nominal over the past 24 hours. No anomalies affecting product delivery or s/c health & safety. For GOES status and other information please see: http://www.oso.noaa.gov/goes/index.htm

POES

0645z Wideband problems with the GE Americom between Wallops and SOCC resulted in very noisy data during the post pass transfers. GE Americom was able to remotely switch the solid state power amps at Wallops and the wideband was returned to service at 0755Z. All data was then reshipped nominally after the wideband link was repaired.

All POES operations were nominal over the past 24 hours No anomalies affecting product delivery or s/c health & safety. For a detailed status and description of POES see: http://www.oso.noaa.gov/poes/index.htm

DMSP

All DMSP operations were nominal over the past 24 hours. No anomalies affecting product delivery or s/c health & safety. For more DMSP Information please see: http://dmsp.ngdc.noaa.gov/dmsp.html

Questions or comments to [email protected]

SpaceRef staff editor.